In tge equation. : 3Fe +4H2O = Fe3O4 + 4H2. if atomic weight of iron is 56 ,then its equivalent weight will be.(what is the valency of Fe in Fe3O4 should be considered.)

Grade:12th pass

5 Answers

Khimraj
7 years ago
In this equivalent weight of Fe is 56/(8/3) = 21. And valency of Fe in Fe3O4 is 4 and 2 as it is mixture of FeO2 and 2 FeO.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
Parul Rajput
7 years ago
How equivalent Weight 8/3 has come how to calculate it .?? ??? Please write the balanced reaction of Fe
Khimraj
7 years ago
As u already wrote ballades equation and as total 8 lection changes so for each Fe atom 8/3 electron changes.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
harsa sannapureddd
5 years ago
Fe3O4 in acidic  medium
is a mixture of Fe2O3 and FeO so equivalent mass is equial to molar mass
 
Fe3O4=fe2o3 + feo
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Fe3O4 in acidic medium
is a mixture of Fe2O3 and FeO so equivalent mass is equial to molar mass
Fe3O4 → Fe2O3 + FeO
n - factor = 8/3
It's equivalent weight = atomic weight/n-factor
= 56×3/8
= 21 gm

Thanks and Regards